# chattersPwCHparticipants with chatbot history

### v1

@@ 6. Conclusion
The evaluation of the GPT language models on the exercises of the SE1 introductory course validates previous findings that current-generation tools are capable of solving the vast majority of, though not all, first-semester programming assignments. For courses where manual reviews of the assignments are integral part—as in our setting—we found that the answer to RQ1 is currently still “no”. Neither GPT-3.5 nor GPT-4 could pass the course without manual intervention, though only by a small margin in the case of GPT-4, due to the strict exercise completion requirements.
@@
In the experimental study, we demonstrated that the use of chatbot assistants is not universal across students. Over a third of our participants chose not to engage with the provided assistant at all. To answer RQ2, we showed that among students who used a chatbot during the task-solving process, two primary strategies emerged: using it as a knowledge base for information retrieval, and using it to generate code solutions. Additionally, we observed ineffective but common patterns of students repeatedly attempting to coax the chatbot into providing a correct solution, by relaying the output of its previous incorrect attempt back to the chatbot.
@@
Most students attempted to solve the problem by themselves first, and did not ask for a code solution in their first message to the chatbot. However, a majority would end up eventually requesting a code solution and trying to submit it. None of the participants asked the model to discuss why their own solution wasn’t behaving as expected. Therefore, to answer RQ3, we found that students do not immediately delegate task-solving to a chatbot, but generally turn to it for a full solution upon encountering difficulties. The willingness to reuse generated code solutions, even if only after failing to solve the problem independently, indicates that concerns around over-reliance on ChatGPT and subsequent academic dishonesty are warranted. Particularly, the tendency of students who report using tools regularly to more strongly favor code generation prompts raises questions about the habits students build through repeated interactions with AI assistants.
@@
In summary, ChatGPT poses a challenge to learning to critically-think in programming courses. If students have unrestricted and unmonitored access to the tool, a significant fraction may use it to avoid autonomous work on tasks they find challenging, without the expertise required to critically evaluate its output and avoid being misled. Provided the course structure allows, shifting focus towards one-on-one reviews of exercises could help uncover cases of AI-related plagiarism, or at least ensure that students actually understand the solutions they are submitting.
More research is needed on the effect of use on learning progress, long-term retention and skill acquisition, as well as the potential of adapting an specifically for an educational context, to determine what place it should have in the classroom moving forward.
7. Data Availibility
@@
The full survey and exercise as well as the interactions and messages recorded from the experiment sessions are available at https://figshare.com/s/d8b532a6ca49b80e6df7. The task-solving evaluation results will be provided before publication.

### v2

@@ 6. Conclusion
The evaluation of the GPT language models on the exercises of our introductory software development course validates previous findings that current-generation widely accessible tools are capable of solving the vast majority of, though not all, first-semester programming assignments. For courses where in-person code reviews of the assignments are an integral part—as in our setting—we found that the answer to RQ1 is “no”, due to practical challenges around justifying and correcting implementation details in AI-generated code.
@@
In the experimental study, we demonstrated that the use of chatbot assistants is not universal across students. Over a third of our participants chose not to engage with the provided assistant at all. Notably, whether the students used the chatbot had no measurable impact on task-solving performance or completion time. To answer RQ2, we showed that among students who used a chatbot during the task-solving process, two primary strategies emerged: using it as a knowledge base for information retrieval, and using it to generate code solutions. We also observed ineffective but common patterns of students repeatedly attempting to coax the chatbot into providing a correct solution, by relaying the output of its previous incorrect attempt back to the chatbot. The students’ continued interactions with the chatbot, despite it consistently generating incorrect solutions, suggests an inability to recognize the limitations of .
@@
Most students attempted to solve the problem by themselves first, and did not ask for a code solution in their first message to the chatbot. However, a majority would end up eventually requesting a code solution and trying to submit it. None of the participants asked the model to discuss why their own solution wasn’t behaving as expected. Therefore, to answer RQ3, we found that students do not immediately delegate task-solving to a chatbot, but generally turn to it for a full solution upon encountering difficulties. The willingness to reuse generated code solutions, even if only after failing to solve the problem independently, indicates that concerns around over-reliance on ChatGPT and subsequent academic dishonesty are warranted. The tendency of students who report using tools regularly to more strongly favor code generation prompts particularly raises questions about the habits students build through repeated interactions with AI assistants.
@@
In summary, ChatGPT poses a challenge to learn programming and develop critically thinking skills. If students have unguided and uncritical access to the tool, a significant fraction may use it to avoid autonomous work on tasks they find challenging, without the expertise required to critically evaluate its output and avoid being misled. Interview-like assessments could help uncover cases of AI-related plagiarism, or at least ensure students actually understand the solutions they are submitting. More research is needed on the effect of on learning, long-term retention, and critically thinking and reasoning skills.
7. Data Availability
@@
The full survey and exercise as well as the interactions and messages recorded from the experiment sessions are available at https://figshare.com/s/d8b532a6ca49b80e6df7.